OpenStudy (anonymous):

what is the slope of (-1,2) (0,4)

OpenStudy (anonymous):

\[\frac{ y-y1 }{ x-x1 }\] That's the equation you use to find slope, if I'm not mistaken. So, plug in the correct numbers.

jigglypuff314 (jigglypuff314):

\[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] given the coordinates (x1, y1) and (x2, y2) so when given (-1 , 2) (0 , 4) plug and solve

OpenStudy (anonymous):

yes i have worked the problem but would like another person to answer to verify its correct.

OpenStudy (texaschic101):

so what did you get ?

OpenStudy (anonymous):

i got 2.

jigglypuff314 (jigglypuff314):

that's correct :)

OpenStudy (texaschic101):

try again..wrong

OpenStudy (anonymous):

yaay thanks!

jigglypuff314 (jigglypuff314):

how would it have been wrong?

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

yes tell me?

jigglypuff314 (jigglypuff314):

4-2 = 2 0- (-1) = 1

OpenStudy (texaschic101):

I am sorry...I made a mistake.....you are correct...it is 2. I was thinking -2. My bad

OpenStudy (anonymous):

ok thanks guys!!